Palencia Paints Corporation has a target capital structure of 45% debt and 55% common equity, with no preferred stock. Its before-tax cost of debt is 13%, and its marginal tax rate is 25%. The current stock price is P0 = $20.50. The last dividend was D0 = $2.00, and it is expected to grow at a 5% constant rate. What is its cost of common equity and its WACC? Do not round intermediate calculations. Round your answers to two decimal places.

rs =   %

WACC =   %
Answer #1

Cost of common equity=(D1/Current price)+Growth rate

=[(2*1.05)/20.5]+0.05

=15.24%(Approx).

Cost of debt after-tax=13*(1-tax rate)

=13*(1-0.25)=9.75%

WACC=Respective costs*Respective weight

=(15.24*0.55)+(0.45*9.75)

=12.77%(Approx).